OpenStudy (anonymous):

X^-2 = (2+or-square root of 3)

OpenStudy (anonymous):

Re-write the negative exponent as a fraction, then take the reciprocal of both sides and raise both sides to the 1/2 power.

OpenStudy (anonymous):

is there any way you could draw it out

OpenStudy (anonymous):

I can get you started.

OpenStudy (anonymous):

thad be great

OpenStudy (anonymous):

\[\large x^{-2} = 2 \pm \sqrt{3}\]\[\large \implies \frac{1}{x^2} = 2\pm \sqrt{3}\]\[\implies x^2 = \frac{1}{2\pm \sqrt{3}}\]\[\implies x^2 = \frac{1}{2\pm \sqrt{3}} \cdot \frac{2 \mp \sqrt{3}}{2\mp \sqrt{3}}\]\[\implies x^2 = \frac{2\pm \sqrt{3}}{4-3} = \frac{2 \pm \sqrt{3}}{1} = 2 \pm \sqrt{3}\]\[\implies x = \pm \sqrt{2 \pm \sqrt{3}}\]

OpenStudy (anonymous):

Actually that's all of it.